40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 00:55:30 2010, Koji, Update, IOO, The dark port beams reached the AP table 
    Reply  Sat Dec 18 01:17:04 2010, kiwamu, Update, IOO, some more alignments 
Message ID: 4067     Entry time: Fri Dec 17 00:55:30 2010     Reply to this: 4070
Author: Koji 
Type: Update 
Category: IOO 
Subject: The dark port beams reached the AP table 

[Koji and Kiwamu]

We obtained two dark port beams on the AP table: OMC REFL and AS

- First, IPANG and BS were aligned so as to have the beams on the center of the ETMs.

- Then ITMX/ITMY/PRM/SRM were aligned to have fringes in a single spot anywhere.

- As we already had the dark port beam on the steering mirrors on the BS table, today the PZT mirrors were adjusted.
This work was the beam steering between the BS table to the OMC table. After some tweaking of the mirror mounts, 
the spot on the last PZT mirror was found.

As we have not touched any of the OMC optics since they were aligned well, the alignment has been adjusted by the nobs of the PZT steering mirrors.
Once the beam is on the output mode matching telescope (OMMT), the work was quite easy thanks to the beam shrinking by the OMMT.

Note that the dark port beam is slightly clipped by the green steering mirror. The steering mirror will be moved next time.

After the alignment, we indeed obtained OMC REFL and AS beams on the AP table.
The fringes were visible on the OMC REFL CCD.

We keep the dark port setup on the OMC (in-vac) and AP tables so that they can be the reference of the dark port alignment.
In principle we can align the beams onto the OMC by the two PZT mirrors.


What is left?

Our minimum success of this vent is to setup the X arm cavity which is needed for the green locking.
This setup was already realized. So we fulfilled the condition to close the tank even if the damping of
the ETMY is not achieved. (But we should try)

Tomorrow, we make a light touches to POY, Green, IPPOS, and check the table leveling, clamping, etc, in general.

JD:  We should check OSEMs for all optics *after* table leveling.  Some of them (esp. BS and ITMX) are currently close to their limits right now.

KA: Check green alignment. / Take photos of the tables. / Fix the leveling weights


Location    Optics            Action
--------------------------------------------------------------
@ITMY -     POY               mirror replacement (45deg->0deg) / alignment

@BSC -      Green steering    alignment
            IPPOS steering    alignment
            Beam dumps
            Table Leveling

@ETMX -     Al foil removal
            Table Leveling

@ETMY -     ETMY damping
            OSEM
            OPLEV
            Table Leveling

@ITM/ETM -  Mirror Wiping

ELOG V3.1.3-